OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). y = -4x + 1 y = -1/2 x + 9 y = 4x + 1 y = -2x + 9

OpenStudy (anonymous):

first we need to find the slope \[slope=\frac{ y2-y1 }{ x2-x1}\] \[\frac{ -1-1 }{ 5-4 }\]

OpenStudy (anonymous):

-2/1 slope

OpenStudy (anonymous):

can that be simplified?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yes it can :) anything whats -2 divided by 1?

OpenStudy (solomonzelman):

you can make a small 2 by x, and by y, say `y_2` or `x_2`

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

now lets put it in \[y-y_1=m(x-x_1)\] using the coordinates (4,1) and because the slope is -2 our m is -2

OpenStudy (anonymous):

-1-1 = -2 (5-4)

OpenStudy (anonymous):

leave the first x and y alone \[y-1=-2(x-4)\]

OpenStudy (anonymous):

can you distribute our slope?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright do that please dont forget 2 negatives will make a positive

OpenStudy (anonymous):

(-2*x) - (-2*4) -2-(-8) -10

OpenStudy (anonymous):

you almost had it right only you would leave -2 alone and add the x and make -8 a positive \[y-1=-2x+8\]

OpenStudy (anonymous):

now to isolate y we have to get -1 on the other side how would we do that?

OpenStudy (anonymous):

add -1 to both sides?

OpenStudy (anonymous):

y = -2x + 9?

OpenStudy (anonymous):

yes you got it :)

OpenStudy (anonymous):

Write an equation of the line that passes through (8, 3) and (4, 3). would this be y=3?